Chair: Lim

Next meeting chair: David Sc

Attended: Simin, David Sc, Ahmad, Jasper, Jay, Jamie, Prad

Apologies: Earl

General Issues: Katherine is still offline, will probably come back on 10-11th April. 1 April trip to Katherine.

Notes:

  • Hb12 now has noise diode and phase cal. Make sure phase cal is on.
  • VGOS test session starting tonight.
  • Jamie will update the documentation once everything is finalised.

Handover issues:

Yg:

  • R41040: fs not communicating with antenna. fs restart did not clear it, so computer needed restart of pcfs machine (DSmith), could be general network issue.
  • R41041: Scheduled halted for restarting dbbc software due to bad spectra (David). It was last two channels, interference maybe.. Progressively worst.
  • aua086: mk5yg needed a restart (Ahmad). Module was the wrong one. mk5yg was not responding. Sometime module swap can lead to this issue, not common.

Hb:

  • aua086: 1818-1854 UT fs not responding (Ahmad) . Kill it didn't do much. Jamie was having a test on new version of fs, had issue reverting it back.